# NavForesee: A Unified Vision-Language World Model for Hierarchical Planning and Dual-Horizon Navigation Prediction

### v1

@@ III-C Model Architecture
Overall Architecture The overall architecture of NavForesee is illustrated in Figure 2. We adopt Qwen2.5-VL-3B-Instruct [36], a large-scale vision–language model with strong multi-modal understanding capabilities, as the backbone. NavForesee is designed to integrate two complementary functionalities: VLM-based language planning and World model-based visual prediction. Correspondingly, we define two primary training objectives: VLM planning training and world model training. Training data from both tasks are jointly mixed to ensure that the model preserves its multi-modal planning ability, while simultaneously extending its capability to generate visual features. For the VLM planning training, textual planning data are directly fed into Qwen for auto-regressive training, leveraging its original text encoder and image encoder components without modification. For the world model training, we introduce an additional position encoder (pos-encoder) to encode the agent’s relative position and orientation from image observations. Two sets of dream queries—corresponding to short- and long-horizon predictions—are appended to the multi-modal embeddings. Each set of dream queries includes depth and semantics subqueries, enabling dual-horizon prediction. Furthermore, an action query, alongside the dream queries, is integrated into the multi-modal inputs and processed by Qwen2.5-VL via a structured attention mechanism. Lightweight convolutional layers serve as decoders to transform dream embeddings into environmental predictions (depth and semantics), while a simple MLP predicts action outputs (waypoints, orientation estimates, and arrival flags).
@@ III-E Predictive Action Policy Learning
where $E_{a}$ is the action embedding and $M_{i}nv$ denotes the inverse dynamics model. Notably, in our action policy learning pipeline, the action embedding $E_{a}$ is extracted through the Qwen2.5-VL backbone, while action predictions are primarily conditioned on the dual-horizon predictive features, ensuring that decisions are informed by both past observations and forecasted environmental dynamics
@@ IV Experimental Evaluation
1 66.2% 78.4% 3.94 59.7%
2 48.8% 75.5% 5.61 39.4%
3 58.6% 76.4% 4.47 50.1%
4 52.6% 67.4% 5.53 46.7%
We evaluate our model in continuous environment of the Habitant simulator on the R2R-CE and RxR-CE datasets.
R2R-CE dataset is derived from the Matterport3D indoor environments, discretized for path planning but operationalized in the Habitat simulator under a continuous navigation setting. It provides fine-grained, step-by-step natural language instructions, allowing for detailed guidance at each navigation step. In the simulator, the embodied agent can execute turns as small as $15^{\circ}$ and perceives the scene through a $90^{\circ}$ horizontal field-of-view.
RxR-CE is a large-scale, multilingual VLN dataset comprising about 126K human-annotated instructions. Compared to R2R-CE, RxR-CE covers more diverse and complex trajectories, increasing the difficulty of the navigation tasks. The agent in this setting uses a coarser minimum turn increment of $30^{\circ}$ and a narrower $79^{\circ}$ horizontal field-ofview, which demands more deliberate movement planning for effective scene coverage.
We evaluate navigation performance using standard metrics including success rate (SR), oracle success rate (OS), success weighted by path length (SPL), and navigation error (NE).
@@ IV-A Comparison with State-of-the-Art Methods
Table 1 reports the performance of the proposed method compared with other approaches on the R2R‑CE and RxR‑CE datasets. Overall, NavForesee delivers competitive results against state‑of‑the‑art (SOTA) methods. Specifically, on the val unseen split of the R2R‑CE dataset, NavForesee achieves SOTA performance by improving SR by 1.1%, OSR by 10.9%, and reducing NE by 0.3 m. This improvement can be attributed to the integration of the world model prediction module, which enables the agent to better capture environmental dynamics, avoid obstacles, and explore the surroundings more effectively.
@@ IV-C Qualitative Analysis
Figure 3 illustrates the short-term depth and semantic feature predictions generated by our world model over the course of a complete navigation episode, forecasting up to four future steps. Although the predicted depth maps appear somewhat coarse—owing to the constraints of pixel-level supervised training on R2R-CE and RxR-CE—they nonetheless preserve the scene’s global geometry and spatial layout, faithfully capturing agent movements such as ascending or descending staircases, entering and exiting rooms, and making sharp or gradual turns. This ability to retain high-level spatial coherence despite reduced pixel detail ensures that the model’s predictions remain informative for downstream navigation decisions. The semantics predictions, obtained via a pretrained segmentation head, exhibit strong alignment with ground truth labels, successfully reflecting dynamic environmental changes in synchrony with the agent’s actions.
@@ I-B Training Details
We interleave the VLM planning training data and world model training data to jointly train NavForesee. The training batch size is set to 4, and the number of image observations is flexible, up to a maximum length of 20. Depth and semantic features are precomputed and loaded during training. We use the AdamW optimizer with an initial learning rate of $1\times 10^{-5}$. Depth and semantics predictions are weighted with $\alpha=0.25$ and $\beta=0.3$. The model is trained for a total of 3 epochs on 64 NVIDIA H20 GPUs, with ViT parameters frozen. The fixed short-term prediction horizon is set to 4, and the number of predicted waypoints is set to 5.
@@ II-B Short-term and Long-term Prediction Evaluation
Figure 6 illustrates the short-term and long-term depth predictions produced by our world model over a complete navigation episode. Short-term predictions forecast up to four future steps, whereas long-term predictions extrapolate over an adaptive horizon determined by progress towards the next milestone. Compared to short-term predictions, long-term depth predictions may be less accurate in capturing detailed depth at milestone locations, since milestone positions are unknown during inference. At the beginning of the episode, the long-term predictions effectively capture the scene when the agent ascends the stairs. As the agent approaches the first milestone (the doorway), the long-term predictions degrade slightly, likely due to the increased uncertainty of longer horizons and the absence of explicit milestone information. In such cases, long-term predictions tend to track short-term outputs, because long-term queries can attend to short-term queries. Nevertheless, the long-term predictions maintain the overall scene layout and depth distribution, providing valuable guidance for strategic navigation. This demonstrates that NavForesee’s world model effectively anticipates environmental changes over both short and long horizons, enhancing the agent’s planning and action capabilities in complex scenarios.
TABLE III: Performance comparison between depth prediction and semantics prediction
@@
1 66.2% 78.4% 3.94 59.7%
2 61.8% 76.7% 4.37 54.9%
3 60.0% 76.2% 4.59 52.9%
II-C Ablation Study on Depth and Semantics Predictions
We conduct ablation studies to evaluate the individual contributions of depth and semantics predictions in the world model. As shown in Table III, removing either depth or semantics predictions results in a clear performance drop. The full NavForesee model, which integrates both depth and semantics predictions, achieves the highest SR (66.2%), OSR (78.4%), lowest NE (3.94), and best SPL (59.7%), validating the benefit of their combination. Without depth prediction, the SR drops to 61.8% and SPL decreases by 4.8 points, highlighting the importance of depth information for spatial reasoning and obstacle avoidance. Disabling semantics predictions leads to an even larger SR reduction (60.0%) and higher NE, underscoring the critical role of semantic features in recognizing landmarks and guiding navigation. These findings confirm that both depth and semantics predictions are essential for accurate and efficient navigation.

### v2

@@ III-C Model Architecture
Overall Architecture The overall architecture of NavForesee is illustrated in Figure 2. We adopt Qwen2.5-VL-3B-Instruct [36] as the backbone. NavForesee is designed to integrate two complementary functionalities: VLM-based hierarchical planning and World model-based visual prediction. Correspondingly, we define two primary training objectives: hierarchical planning training and world model training. We optimize the hierarchical planning and world model objectives in a unified manner by interleaving their respective training data, ensuring that the model preserves its multi-modal planning ability while simultaneously extending its capability to generate visual features. For the hierarchical planning training, textual planning data are directly fed into Qwen for auto-regressive training, leveraging its original encoder components without modification. For the world model training, we introduce a position encoder and integrate dream queries (covering short/long-horizon depth and semantics) and action queries into the multi-modal input. These are processed via structured attention, where lightweight decoders transform dream embeddings into environmental features and an MLP predicts navigation actions. More details are in the supplementary material.
@@ III-D Dual-horizon World Model Prediction
The long-term prediction horizon $M_{t}$ is not rigid or preset. While sub-instructions are learned implicitly during hierarchical planning, the planning-aware hidden states effectively encode the current sub-goal intent. This alignment is achieved through the interleaved training of planning and prediction tasks via shared representations.
@@ III-E Predictive Action Policy Learning
where $E_{a}$ is the action embedding and $M_{inv}$ denotes the inverse dynamics model. Notably, in our action policy learning pipeline, the action embedding $E_{a}$ is extracted through the Qwen2.5-VL backbone, while action predictions are primarily conditioned on the dual-horizon predictive features, ensuring that decisions are informed by both past observations and forecasted environmental dynamics.
@@ IV Experimental Evaluation
1 66.2 78.4 3.94 59.7
2
48.8 75.5 5.61 39.4
3
47.8 75.3 5.77 36.1
4
58.6 76.4 4.47 50.1
5
52.6 67.4 5.53 46.7
We evaluate our model using the Habitat simulator on the R2R-CE and RxR-CE datasets.
R2R-CE is based on Matterport3D environments and provides fine-grained, step-by-step instructions, allowing for detailed guidance at each navigation step. In the simulator, the embodied agent can execute turns as small as $15^{\circ}$ and perceives the scene through a $90^{\circ}$ horizontal field-of-view.
RxR-CE is a large-scale, multilingual VLN dataset comprising about 126K human-annotated instructions featuring more complex trajectories. The agent in this setting uses a coarser minimum turn increment of $30^{\circ}$ and a narrower $79^{\circ}$ horizontal field-ofview, which demands more deliberate movement planning for effective scene coverage.
We report standard metrics: Success Rate (SR), Oracle Success Rate (OS), Success weighted by Path Length (SPL), and Navigation Error (NE).
Training details are provided in the supplementary material.
@@ IV-A Comparison with State-of-the-Art Methods
Table I reports the performance of the proposed method compared with other approaches on the R2R‑CE and RxR‑CE datasets. Note that methods marked with an asterisk (*) denote discrete waypoint-based approaches. Overall, NavForesee demonstrates superior performance compared to state-of-the-art (SOTA) methods.
Specifically, on the val unseen split of the R2R‑CE dataset, NavForesee achieves SOTA performance by improving SR by 1.1%, OSR by 10.9%, and reducing NE by 0.3 m. This improvement can be attributed to the integration of the world model prediction module, which enables the agent to better capture environmental dynamics, avoid obstacles, and explore the surroundings more effectively.
@@ IV-B Ablation Study
We vary the short-term prediction horizon ($k$) from 3 to 5 (waypoint length) to validate the choice of short-term horizon. Results in Table III confirm that matching the horizon to the action space ($k=5$) is optimal.
TABLE III: Ablation study of Short-term horizon.
short-term horizon SR $\uparrow$ OSR $\uparrow$ NE $\downarrow$ SPL $\uparrow$
k=5 66.2 78.4 3.94 59.7
k=4 64.1 78.6 4.08 54.9
k=3 56.5 77.2 4.85 48.5
@@ IV-C Qualitative Analysis
Figure 3 illustrates the short-term depth and semantic feature predictions generated by our world model over the course of a complete navigation episode, forecasting up to five future steps. Although the predicted depth maps appear somewhat coarse—owing to the constraints of pixel-level supervised training on R2R-CE and RxR-CE—they nonetheless preserve the scene’s global geometry and spatial layout, faithfully capturing agent movements such as ascending or descending staircases, entering and exiting rooms, and making sharp or gradual turns. This ability to retain high-level spatial coherence despite reduced pixel detail ensures that the model’s predictions remain informative for downstream navigation decisions. The semantics predictions, obtained via a pretrained segmentation head, exhibit strong alignment with ground truth labels, successfully reflecting dynamic environmental changes in synchrony with the agent’s actions.
@@ I-B Training Details
We interleave the VLM planning training data and world model training data to jointly train NavForesee. The training batch size is set to 4, and the number of image observations is flexible, up to a maximum length of 20. Depth and semantic features are precomputed and loaded during training. We use the AdamW optimizer with an initial learning rate of $1\times 10^{-5}$. Depth and semantics predictions are weighted with $\alpha=0.25$ and $\beta=0.3$. The model is trained for a total of 3 epochs on 64 NVIDIA H20 GPUs, with ViT parameters frozen. The fixed short-term prediction horizon is set to 5, same as the number of predicted waypoints.
@@ II-B Short-term and Long-term Prediction Evaluation
Figure 6 illustrates the short-term and long-term depth predictions produced by our world model over a complete navigation episode. Short-term predictions forecast up to five future steps, whereas long-term predictions extrapolate over an adaptive horizon determined by progress towards the next milestone. Compared to short-term predictions, long-term depth predictions may be less accurate in capturing detailed depth at milestone locations, since milestone positions are unknown during inference. At the beginning of the episode, the long-term predictions effectively capture the scene when the agent ascends the stairs. As the agent approaches the first milestone (the doorway), the long-term predictions degrade slightly, likely due to the increased uncertainty of longer horizons and the absence of explicit milestone information. In such cases, long-term predictions tend to track short-term outputs, because long-term queries can attend to short-term queries. Nevertheless, the long-term predictions maintain the overall scene layout and depth distribution, providing valuable guidance for strategic navigation. This demonstrates that NavForesee’s world model effectively anticipates environmental changes over both short and long horizons, enhancing the agent’s planning and action capabilities in complex scenarios.
II-C Quantitative prediction Quality
We evaluate prediction quality on R2R Val-Unseen (input at $T$, target at T+5T{+}5). To verify that the model learns temporal dynamics rather than copying inputs, we compare against Identity Baselines that use GT at T+iT{+}i as predictions for T+5T{+}5. Table IV shows: 1) Depth: Our model outperforms baselines up to T+2T{+}2, capturing short-term geometric dynamics. 2) Semantics (DINO): Our model achieves a CosSim of 0.62, significantly outperforming baselines up to T+3T{+}3.
This indicates strong capability in predicting high-level semantic evolution further into the future.
TABLE IV: Prediction Quality. Model (Pred) vs. Identity Baselines (copying GT at T+iT{+}i) as predictions for T+5T{+}5.
Type Metric Model Identity Baselines (GT vs T+5)
Pred $T$ T+1T{+}1 T+2T{+}2 T+3T{+}3 T+4T{+}4
Depth SSIM $\uparrow$ 0.81 0.79 0.80 0.82 0.84 0.88
PSNR $\uparrow$ 16.39 15.03 15.39 15.98 17.00 19.04
LPIPS $\downarrow$ 0.28 0.32 0.31 0.28 0.25 0.18
Semantic(DINO) CosSim $\uparrow$ 0.62 0.47 0.50 0.53 0.58 0.67
II-D Ablation Study on Depth and Semantics Predictions
We conduct ablation studies to evaluate the individual contributions of depth and semantics predictions in the world model. As shown in Table V, removing either depth or semantics predictions results in a clear performance drop. The full NavForesee model, which integrates both depth and semantics predictions, achieves the highest SR (66.2%), OSR (78.4%), lowest NE (3.94), and best SPL (59.7%), validating the benefit of their combination. Without depth prediction, the SR drops to 61.8% and SPL decreases by 4.8 points, highlighting the importance of depth information for spatial reasoning and obstacle avoidance. Disabling semantics predictions leads to an even larger SR reduction (60.0%) and higher NE, underscoring the critical role of semantic features in recognizing landmarks and guiding navigation. These findings confirm that both depth and semantics predictions are essential for accurate and efficient navigation.
Comparing Table IV with Table V, we conclude that high-quality semantic foresight is more critical for navigation performance than short-term geometric precision, as it aids in long-term goal anchoring.
TABLE V: Performance comparison between depth prediction and semantics prediction
@@ II-D Ablation Study on Depth and Semantics Predictions
1 66.2 78.4 3.94 59.7
2
61.8 76.7 4.37 54.9
3
60.0 76.2 4.59 52.9